OpenStudy (anonymous):

Given the function f(x) = 3^(x), find the value of f^(-1)(81).

OpenStudy (anonymous):

We need the inverse again. Remember how?

OpenStudy (anonymous):

y = 3^x x = 3^y Solve for y from here.

OpenStudy (anonymous):

What are your answer choices I am getting something weird

OpenStudy (anonymous):

f^(-1)(81) = 1 f^(-1)(81) = 2 f^(-1)(81) = 3 f^(-1)(81) = 4

OpenStudy (anonymous):

so would it just be C?

OpenStudy (anonymous):

I get inverse = \[\frac{ \ln (x) }{ \ln (3) }\]

OpenStudy (anonymous):

\[\frac{ \ln (81) }{ \ln (3) }\]

OpenStudy (anonymous):

i'm kind of confused as to how you got that

OpenStudy (anonymous):

hold on one second

OpenStudy (anonymous):

okie dokie

OpenStudy (anonymous):

to be honest, I just checked it on an inverse calculator and that's what i got, so

OpenStudy (anonymous):

I think the answer will be 4. I put in ln81 and got 4.39444915467 and ln3 and got 1.09861228867 Divide them: 4.39444915467/1.09861228867 = 3.999999 So 4

OpenStudy (anonymous):

Hope that is ok...

OpenStudy (anonymous):

ah okay i see what you did now

OpenStudy (anonymous):

ok cool